### v1

# Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal

Abstract
During the Cosmic Dawn (CD), the HI 21-cm optical depth ($\tau$ ) in the intergalactic medium can become significantly large. Consequently, the second and higher-order terms of $\tau$ appearing in the Taylor expansion of the HI 21-cm differential brightness temperature ($\delta T_{\rm b}$ ) become important. This introduces additional non-Gaussianity into the signal. We study the impact of large $\tau$ on statistical quantities of HI 21-cm signal using a suite of standard numerical simulations that vary X-ray heating efficiency and the minimum halo mass required to host radiation sources. We find that the higher order terms suppress statistical quantities such as skewness, power-spectrum and bispectrum. However, the effect is found to be particularly strong on the non-Gaussian signal. We find that the change in skewness can reach several hundred percent in low X-ray heating scenarios. For moderate and high X-ray heating models changes are around $\sim 40\%$ and $60\%$, respectively, for $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. This change is around $\sim 75\%$, $25\%$ and $20\%$ for low, moderate and high X-ray heating models, respectively, for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$. The change in bispectrum in both the halo cutoff mass scenarios ranges from $\sim 10\%$ to $\sim 300\%$ for low X-ray heating model. However, for moderate and high X-ray heating models the change remains between $\sim 10\%$ to $\sim 200\%$ for both equilateral and squeezed limit triangle configuration. Finally, we find that up to third orders of $\tau$ need to be retained to accurately model $\delta T_{\rm b}$, especially for capturing the non-Gaussian features in the Hi 21-cm signal.
keywords:
radiative transfer - galaxies: formation - intergalactic medium - cosmology: theory - dark ages, reionization, first stars - X-rays: galaxies
pubyear: 2025pagerange: Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian SignalEffects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal
1 Introduction
Observations of the cosmological Hi 21-cm signal hold great promise for unveiling the details of the Cosmic Dawn (CD) and the Epoch of Reionization (EoR) (Pritchard & Loeb, 2012; Bera et al., 2023). Significant efforts are currently underway to study this era through radio observations of the Hi 21-cm line. Ongoing interferometric telescopes such as the Low Frequency Array (LOFAR [^1] http://www.lofar.org/; van Haarlem et al., 2013), the upgraded Giant Metrewave Radio Telescope (uGMRT; Ali et al., 2008; Ghosh et al., 2012; Shaw et al., 2023), the Murchison Widefield Array (MWA [^2] https://www.mwatelescope.org/telescope/; Tingay et al., 2013), the Hydrogen Epoch of Reionization Array (HERA [^3] https://reionization.org/; DeBoer et al., 2017) and the New Extension in Nançay Upgrading LOFAR (NenuFAR [^4] https://nenufar.obs-nancay.fr/en/homepage-en/; Munshi et al., 2024) along with the upcoming Square Kilometre Array (SKA [^5] http://www.skatelescope.org/; Mellema et al., 2013; Koopmans et al., 2015), have the primary aim of measuring statistical properties of the signal, such as the power spectrum and bispectrum, to probe the early universe. Some upper limits on the 21-cm power spectrum have already been reported using observations with LOFAR (Mertens et al., 2020, 2025; Acharya et al., 2024), GMRT (Paciga et al., 2013; Pal et al., 2021), PAPER (Cheng et al., 2018; Kolopanis et al., 2019), MWA (Barry et al., 2019; Trott et al., 2020) and HERA (Abdurashidova et al., 2023) and are being used to constrain some of the source and IGM properties during CD-EoR (see e.g. Mondal et al., 2020; Ghara et al., 2020, 2021; Greig et al., 2021; Abdurashidova et al., 2022; Ghara et al., 2025). In this context, accurately modelling the statistical properties of the signal is crucial. It not only deepens our understanding of the signal itself but also helps in designing effective observational strategies, data analysis methods, and ensuring reliable interpretation of any detection.
In this paper, we study the impact of large 21-cm signal optical depth on the spatial fluctuations of the CD 21-cm signal differential brightness temperature ($\delta T_{\rm b}$). In particular, we focus on the non-Gaussian features of the CD 21-cm signal, along with its power spectrum. The 21-cm signal optical depth ($\tau$), which is proportional to the Hi density and inversely proportional to the Hi spin temperature, can become significantly large during CD, particularly in regions with high Hi density and low spin temperature. As a result, when $\tau$ is expanded in Taylor series, not only the linear term (which has been generally taken into account in the literature), but also contributions from the second and higher-order terms in the expansion of $\tau$ to the $\delta T_{\rm b}$ equation become important (see e.g., Datta et al., 2022). These higher-order terms not only change the spatial fluctuations in $\delta T_{\rm b}$, but also introduce additional non-Gaussianity in the signal. Therefore, it is important to investigate how this affects the spatial fluctuations in $\delta T_{\rm b}$, as well as the non-Gaussian properties of the signal from CD.
There have been earlier studies of the non-Gaussianity in the Hi 21-cm signal from CD and reionization eras using statistical measures such as skewness and the skew spectrum (Harker et al., 2009; Patil et al., 2014; Ross et al., 2021). These statistical quantities have been shown to be detectable by the SKA-Low (Ma & Peng, 2023; Cook et al., 2024).
The bispectrum has been extensively used to quantify the non-Gaussian nature of the Hi 21-cm signal from the CD-EoR (Bharadwaj & Pandey, 2005; Shimabukuro et al., 2016; Majumdar et al., 2018). The bispectrum, a Fourier conjugate of the three-point correlation function, is sensitive to correlations between different Fourier modes. Majumdar et al. (2020) presented the first comprehensive analysis of the EoR 21-cm bispectrum across all unique $k$-triangle configurations. Subsequent studies have explored various physical effects on the 21-cm bispectrum, including X-ray heating and $\rm{Ly{\alpha}}$ coupling (Watkinson et al., 2018; Kamran et al., 2021b, a, 2022), IGM physics (Kamran et al., 2021a, 2022), redshift-space distortions and light cone effect (Kamran et al., 2021b; Mondal et al., 2021; Gill et al., 2024), impact of ionizing source models and ionization morphology (Noble et al., 2024; Hutter et al., 2019), impact of distribution of neutral hydrogen islands at the end stage of reionization (Raste et al., 2024), impact of different kinds of dark matter models (Saxena et al., 2020) etc. It has also been demonstrated that the SKA-Low will be sensitive enough to detect the bispectrum signal from the reionization era (Mondal et al., 2021; Tiwari et al., 2022) and bispectrum can put tighter constraints on reionization model parameters (Tiwari et al., 2022) compared to the signal power spectrum. Furthermore, a fast and scalable estimator has been developed for computing the bispectrum of the redshifted 21 cm signal from radio interferometric observations (Gill & Bharadwaj, 2025). Efforts have also been made to directly measure the bispectrum of the signal from observational data (Trott et al., 2019).
Given the promising prospects of detecting the CD-EoR 21-cm power spectrum and bispectrum with ongoing and upcoming radio interferometric experiments, and their potential to constrain CD-EoR models, we present a detailed study on the impact of higher-order terms of the Hi optical depth on the spatial fluctuations of $\delta T_{\rm b}$, as well as its non-Gaussian features. In our earlier work (Datta et al., 2022), we investigate the impact of large Hi 21-cm optical depth on the interpretation of measured global signal. Further, we carried out a limited study on the impact of optical depth corrections on various statistical measures of the $\delta T_{\rm b}$ signal, such as its variance, skewness, and power spectrum, at a few redshifts during CD. Here, our analysis focuses specifically on quantifying non-Gaussianity in the signal through skewness and the bispectrum. We use a suite of standard numerical simulations that vary key parameters such as the X-ray heating efficiency and the minimum halo mass required to host radiation sources. These variations result in a broad range of Hi 21-cm optical depths, enabling us to examine the effects across a wide variety of possible scenarios throughout the CD. In addition, we investigate in detail the impact of optical depth corrections on the Hi power spectrum, which remains the primary target of most current experimental efforts.
The outline of the paper is as follows. In Section 2 we will discuss about the fundamentals of the Hi 21-cm signal. We present the basic equations for computing the Hi 21-cm differential brightness temperature, both exactly and approximately as outlined in subsection 2.1. Subsection 2.2 describes the simulations and models used in this work along with the results on the global evolution of the Hi 21-cm signal. In Section 3 we discuss our results. Subsections 3.1 and 3.2 discusses the optical depth, and its distribution and the impact of large Hi optical depth on the non-Gaussian features of the 21-cm signal, focusing on the distribution of $\delta T_{\rm b}$. In subsections 3.3 and 3.4, we discussed the redshift evolution of the skewness and the impact of large optical depth on it, respectively. Subsections 3.5 and 3.6 examine the redshift evolution of the bispectrum and explores how large optical depths influence it. We also discussed the effect of large optical depth on the Hi 21-cm power spectrum in Section 4. Section 5 investigates the effect of including different orders in the optical depth expansion on bispectrum estimation. Finally, Section 6 summarizes the main results of the paper.
In this paper, we use the values of various cosmological parameters for our numerical calculations and simulations: $\Omega_{\rm b}=0.044$, $\Omega_{\rm m}=0.27$, $h=0.7$ and the present CMB temperature of $2.725$ K. These values are consistent with the WMAP results (Hinshaw et al., 2013) and lie within the error bars of the Planck results (Ade et al., 2014).
2 21-cm signal from CD & EoR
In this section, we present the basic equations governing the Hi $21$-cm signal, along with a detailed description of the simulations we use in this study.
2.1 Basic equations of Hi  21-cm signal
The differential brightness temperature corresponding to redshifted Hi $21$-cm signal relative to the CMBR can be written as (Rybicki & Lightman, 1986),
$\delta T_{\rm b}({\bf n},z)=\frac{T_{S}({\bf n},z)-T_{\gamma}(z)}{1+z}\left[1-\exp\{-\tau({\bf n},z)\}\right],$ (1)
where $T_{S}({\bf n},z)$ is the Hi spin temperature corresponding to its hyperfine transition and $T_{\gamma}$ is the CMBR temperature. $\tau({\bf n},z)$ is the optical depth that quantifies the level of absorption or emission of Hi 21-cm line within a medium (here IGM) and can be written as (Pritchard & Loeb, 2008),
$\tau({\bf n},z)$ $=$ $27\hskip 2.84544pt\frac{x_{\rm HI}(\textbf{n},z)}{T_{S}(\textbf{n},z)}\left[1+\delta_{\rm b}(\textbf{n},z)\right]\left(\frac{\Omega_{\rm b}h^{2}}{0.023}\right)$ (2)
$\times$ $\left(\frac{0.15}{\Omega_{\rm m}h^{2}\cdot 10}\right)^{1/2}(1+z)^{3/2}\,\mathrm{mK},$
where $x_{\rm HI}(\textbf{n},z)$ and $\delta_{\rm b}(\textbf{n},z)$ are neutral hydrogen fraction and baryonic density contrast respectively along the line of sight direction n and redshift $z$. $\Omega_{\rm b}$ and $\Omega_{\rm m}$ are the cosmological baryon density and total matter density parameters at present, respectively. $h$ is the present Hubble constant in units of $100$ km/s/Mpc. We do not consider the effects of the peculiar velocity in this study.
Normally, it is assumed that $\tau({\bf n},z)\ll 1$, and under this assumption the equation (1) can be approximated as,
$\delta T_{\rm b}({\bf n},z)=\frac{T_{S}({\bf n},z)-T_{\gamma}(z)}{1+z}\tau({\bf n},z),$ (3)
which can be further written as,
$\delta T_{\rm b}(\textbf{n},z)=27x_{\rm HI}(\textbf{n},z)\big{[}1+\delta_{\rm b}(\textbf{n},z)\big{]}\left(\frac{\Omega_{\rm b}h^{2}}{0.023}\right)$
×$\times\left(\frac{0.15}{\Omega_{\rm m}h^{2}}\frac{1+z}{10}\right)^{1/2}\left[1-\frac{T_{\gamma}(z)}{T_{\rm S}(\textbf{n},z)}\right]\,\mathrm{mK}.$ (4)
The above equation is widely used for modelling the signal, both analytically and numerically. However, the optical depth can be very large at high redshifts in places where the spin temperature is low and the Hi density is large. During CD, the spin temperature can be very low before the onset of X-ray heating when it is coupled to the IGM kinetic temperature. This leads to a large Hi optical depth, particularly in regions with high Hi density. For larger values of $\tau$, it is necessary to include higher-order terms, such as $\tau^{2}$, $\tau^{3}$, and beyond, in equation (3) to accurately compute $\delta T_{\rm b}(\textbf{n},z)$. However, for moderate $\tau$ values equation (1) can be approximated as,
$\delta T_{\rm b}(\textbf{n},z)=\frac{T_{S}(\textbf{n},z)-T_{\gamma}(z)}{1+z}\left[\tau-\tau^{2}/2+\tau^{3}/6-\mathcal{O}(\tau^{4})\right].$ (5)
This inclusion of higher-order terms would influence various statistics of the signal.
2.2 Simulations of Hi  21-cm signal during CD
We use grizzly code (Ghara et al., 2015a, b, 2018) for $21$-cm signal simulations, which is based on a one-dimensional radiative transfer scheme. This code requires a gas density field, a halo catalog and a source model as its inputs. The dark matter distribution in a cubic box of comoving length $714.29$ Mpc is obtained using an N-body code named cubep3m (Harnois-Deraps et al., 2013) which is smoothed into 600 uniform grids resulting in a spatial resolution of $1.19$ Mpc. Dark matter distributions are obtained at multiple redshifts in the range of $z=18$ to $z=10$ during CD. Dark matter halos are identified using an on-the-fly halo finder which uses the spherical overdensity method. The lowest dark matter halos identified in our simulations have masses of approximately $10^{9}$ ${\rm M}_{\odot}$ (the mass of a dark matter particle in our simulations is $\sim 10^{8}$ ${\rm M}_{\odot}$). The stellar mass available within a halo is directly related to the DM halo mass as $M_{*}=f_{*}(\frac{\Omega_{\rm b}}{\Omega_{\rm m}})M_{\rm h}$, where the star formation efficiency $f_{*}$ is assumed to be $0.03$ (Behroozi & Silk, 2015; Sun & Furlanetto, 2016). Our simulation has two different astrophysical parameters i.e., $f_{\rm x}$ and $M_{\rm h,min}$. $M_{\rm h,min}$ is the minimun mass of a dark matter halo capable of hosting sources sources that emit $\rm{Ly{\alpha}}$, X-ray and UV ionizing photons. We consider two different values of $M_{\rm h,min}$, i.e. $10^{9}$ ${\rm M}_{\odot}$ and $10^{10}$ ${\rm M}_{\odot}$, in this study. $f_{\rm x}$ is the X-ray heating efficiency parameter which is the ratio of X-ray to UV luminosity of a source. Table 1 summarizes the model parameters considered in this study. The details of these simulations are discussed in (Kamran et al., 2022).
Using the DM density distribution obtained from the N-body simulation and spin temperature, neutral hydrogen distribution from the one-dimensional radiative transfer code grizzly, we generate the Hi 21-cm optical depth $\tau({\bf n},z)$ distribution and subsequently the Hi $21$-cm differential brightness temperature ($\delta T_{b}({\bf n},z)$) maps. We generate $\delta T_{b}$ maps at $23$ different redshifts covering the entire CD using both the approximated (equation (3)) and the exact (equation (1)) $\delta T_{\rm b}$ equations.
We show the evolution of the global signal ($\overline{\delta T_{\rm b}}(z)$) with redshift in the upper panel of Fig. 1. We see, for all choices of $f_{\rm x}$ values and for a fixed $M_{\rm h,min}$, that the initial trend in the evolution of $\overline{\delta T_{\rm b}}(z)$ around redshift $z=18$ is very similar as only $\rm{Ly{\alpha}}$ coupling dictates the 21-cm signal during the initial phase. It helps to couple the spin temperature with the IGM kinetic temperature; therefore, the signal is seen in absorption. Later, X-ray heating becomes important and raises the IGM kinetic and spin temperature. This creates a trough-like feature in the redshift evolution of $\overline{\delta T_{\rm b}}(z)$. Further, we see that the absorption trough becomes deeper for smaller $f_{\rm x}$ values as the IGM heating starts late and the spin temperature, which is coupled to IGM kinetic temperature through $\rm{Ly{\alpha}}$ coupling, continues to decrease due to adiabatic expansion of universe. Moreover, the trough is more flattened for smaller $f_{\rm x}$ values. This is because the X-ray heating rate and cooling rate due to cosmological expansion remain similar for a longer period of time in these scenarios. As expected, higher $f_{\rm x}$ makes $\overline{\delta T_{\rm b}}(z)$ positive at higher redshift. The trough like feature in $\overline{\delta T_{\rm b}}(z)$ shifts towards lower redshifts for higher halo mass cut-off, mainly because massive DM halos are formed later and, therefore, processes such as $\rm{Ly{\alpha}}$ coupling, X-ray heating get delayed.
$f_{\rm X}$ $0.1$ $46.4$ $1000$
$M_{\rm h,min}\,(10^{9}\,{\rm M}_{\odot})$ $1.0$ $10.0$
Table 1: Different model parameters of X-ray heating efficiency $f_{\rm X}$ and minimum halo mass $M_{\rm h,min}$.
Figure 1: Redshift evolution of the mean HI 21-cm optical depth and the differential brightness temperatures. The upper and lower panels show the redshift evolution of mean Hi 21-cm optical depth, $\tau$, and the global 21-cm differential brightness temperature $\overline{\delta T_{\rm b}}(z)$, respectively. In the lower panel, the evolution of both approximated (solid curve) and exact $\overline{\delta T_{\rm b}}(z)$ (dashed curve) is shown.
From the lower panel we also see that if we vary the X-ray heating efficiency $f_{\rm x}$, both the width and depth of the troughs vary. These are maximum for $f_{\rm x}=0.1$ and minimum for $f_{\rm x}=1000$. Further, we see that $\overline{\delta T_{\rm b}}(z)$ calculated using the exact equation is smaller than the approximated method, and the difference is more for smaller values of $f_{\rm x}$. This is expected as $T_{\rm S}$ has lower values for the $f_{\rm x}=0.1$ model, i.e., $\tau$ has higher values, which leads to a greater difference in the troughs between approximated and exact $\overline{\delta T_{\rm b}}(z)$ compared to the $f_{\rm x}=1000$ model. The difference is found to be similar for different $M_{\rm h,min}$ models for a given $f_{\rm x}$ value. The percentage difference between approximated and exact $\overline{\delta T_{\rm b}}(z)$ reaches up to $6.2\%$ for [$f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$] model and $6.7\%$ for [$f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$] model. The differences reduce to roughly $5\%$ and $2.7\%$ for $f_{\rm x}=46.4$ and $1000$, respectively, for both the $M_{\rm h,min}$ values. This difference arises because of considerably large values of $\tau$ during the CD. The higher-order terms in $\exp(-\tau)$ become significant when $\tau$ is large. To elaborate on this, we plot the evolution of the mean $\tau$ with redshift in the upper panel of Fig. 1. The redshift at which the mean $\tau$ reaches its maximum corresponds to the largest difference between the exact and approximated $\overline{\delta T_{\rm b}}(z)$. The optical depth becomes maximum when the IGM kinetic and spin temperature are minimum, which occurs just before the onset of X-ray heating.
3 Results
This section presents results on the distributions of HI 21-cm optical depth and differential brightness temperature. It further shows results on the redshift evolution of skewness and bispectrum, as well as the impact of large optical depth on these quantities.
3.1 Hi  21-cm optical depth ($\tau$) distribution
We plot the histogram and the cumulative distribution function (CDF) of simulated $\tau$ at redshift $z=12.3$ for three models in Fig. 2. The average $\tau$ values at $z=12.3$ for these models are higher compared to other redshifts, which can be seen in the upper panel of Fig. 1.
Figure 2: This shows histograms of Hi 21-cm optical depth ($\tau$) for models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$ (top left); $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ (top right) and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ (bottom left), at redshift $z=12.3$. The bottom right panel shows the cumulative distribution function (CDF) of $\tau$ for all three scenarios.
We see from the CDF (refer to the bottom right panel of Fig.2) that as many as $\sim 55\%$ of pixels in the simulated cube have $\tau$ value greater than $0.1$ for the model corresponding to $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$. Even for higher X-ray efficiency models corresponding to $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$, around $49\%$ and $25\%$ of the total pixels- respectively, where spin temperature is relatively higher- have $\tau$ value greater than $0.1$. Furthermore, we observe that in all scenarios, the distribution of $\tau$ is positively skewed. The optical depth ($\tau$) depends on both the matter density and the spin temperature (refer to equation (2)), both of which are non-Gaussian. This results in a non-Gaussian $\tau$ field, and consequently contributes to a non-Gaussian $\delta T_{b}$ field during CD.
3.2 Non-Gaussian Hi 21-cm signal
Fig.3 shows the probability distribution functions (PDF) of $\delta T_{\rm b}$ estimated using the approximated and exact equations for three different scenarios of CD. For these models we find that $\delta T_{\rm b}$ distribution is non-Gaussian and negatively skewed (Fig.4) at $z\approx 12.3$. We further see that the exact $\delta T_{\rm b}$ is less skewed compared to the approximated one. The reason is that, as revealed in our simulations, $\tau$ can become considerably large during the Cosmic Dawn. Therefore, higher-order terms like $\tau^{2}$ and $\tau^{3}$ in $\delta T_{b}$ (refer to eq. 5) can no longer be neglected. An accurate calculation of $\delta T_{\rm b}$ reduces its skewness compared to the approximated calculation due to the presence of higher-order terms of $\tau$ in the exact equation. We further notice that $\delta T_{\rm b}$ with large negative values are more suppressed, as these $\delta T_{\rm b}$s are associated with large $\tau$ values, and therefore, are more affected. Voxels with small $|\delta T_{\rm b}|$ values are less affected as these are associated with very small $\tau$. Therefore, large optical depth not only suppresses the overall fluctuations in $\delta T_{b}$ but also impacts the non-Gaussian nature of the signal. In other words, we can say that higher-order terms such as $\tau^{2}$, $\tau^{3}$ introduce additional non-Gaussianity to $\delta T_{\rm b}$ distribution.
Figure 3: This shows the probability distribution function (PDF) of simulated $\delta T_{\rm b}$ for models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$; $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$, at redshift $z=12.3$.
3.3 Skewness and its redshift evolution
Figure 4: This shows the redshift evolution of skewness for the approximated and exact $\delta T_{\rm b}$ (left panels) and the percentage change (right panels) for all the CD scenarios. In the left panel the blue solid and red dashed lines represent the skewness for the approximated and exact $\delta T_{\rm b}$ respectively.
The presence of a large numbers of $\rm{Ly{\alpha}}$-coupled regions during the early CD, heated regions during the heating phase, and ionized bubbles during reionization enhances the non-Gaussian signal and its skewness. As $\rm{Ly{\alpha}}$ coupling and IGM heating progress toward completion, non-Gaussianity gradually decreases. $\delta T_{\rm b}$ being positive inside heated regions makes the distribution positively skewed. On the other hand, it is negative inside the $\rm{Ly{\alpha}}$ coupled regions before heating, resulting in a negatively skewed signal.
In order to quantify non-Gaussianity in the signal and impact of higher-order terms of $\tau$, we estimate the skewness of $\delta T_{\rm b}$ as a function of redshift for different scenarios. Skewness is a measure of asymmetry in the PDF from a Gaussian distribution, and the simplest one-point statistic that is sensitive to non-Gaussianity. Skewness for $\delta T_{\rm b}$ is defined as,
$\rm{skewness}=\left\langle\left\{\frac{\delta T_{\rm b}({\bf n},z)-{\overline{\rm\delta T}}_{b}(z)}{\sigma_{\delta T_{\rm b}}}\right\}^{3}\right\rangle,$ (6)
where $\sigma_{\delta T_{\rm b}}$ is the standard deviation of the $\delta T_{\rm b}$ field.
The left column of Fig.4 shows skewness as a function of redshift $z$ for all scenarios considered here. We can see that the skewness is negative from the beginning of CD until the mean IGM temperature remains lower than the CMBR temperature. Furthermore, the skewness calculated using the accurate $\delta T_{b}$ is smaller in magnitude as compared to the approximated one. The right column of Fig. 4 shows the percentage difference between the two estimates, which quantifies the effect of higher-order terms. We find that all models show significant deviations in skewness. These deviations tend to occur at the redshifts where $\tau$ is large.
At the beginning of CD, skewness is largely negative around redshift $z=18$. As redshift decreases, its magnitude gradually reduces until the heating of the IGM begins. However, during the initial stages of X-ray heating the skewness starts becoming more negative again. As redshift decreases further, heated regions around sources grow significantly and percolates through the IGM. As a result, skewness becomes positive and continues to increase. Eventually, when the entire IGM is nearly heated, the skewness decreases again. For example, in the model with $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ and $f_{\rm x}=1000$, skewness is $\approx-30$ at redshift $z\approx 18$. Its magnitude gradually decreases until redshift $z\sim 14$ as the $\rm{Ly{\alpha}}$ coupling spreads through the IGM and nears completion. Following this, X-ray heating starts, and skewness tends towards more negative values until redshift $z\approx 12.5$. However, as the heated regions expand and percolates through the IGM, the skewness quickly becomes positive, reaching a peak around $z\approx 11.5$. Later, skewness decreases again as almost the entire IGM gets heated. A similar pattern is observed in other scenarios. However, the actual timings of these processes differ depending on the values of $M_{\rm h,min}$ and $f_{\rm x}$. In general, higher values of $M_{\rm h,min}$ and/or lower values of $f_{\rm x}$ delay these processes.
We can further see from the left column of Fig. 4 that the initial trend in the redshift evolution of skewness is almost similar for a fixed $M_{\rm h,min}$ value when $\rm{Ly{\alpha}}$ coupling dominates. This redshift range is from the beginning up to $z\approx 16.5$ for $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$ and upto $z\sim 14$ for $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ models. For larger value of $f_{\rm x}$, X-ray heating starts earlier and dominates the $\rm{Ly{\alpha}}$ coupling, shortening its redshift window (lower panel of Fig. 1).
Based on the results above, we conclude that large heated and Ly-$\alpha$ coupled regions enhance the non-Gaussianity of the signal. The completion of $\rm{Ly{\alpha}}$ coupling and IGM heating, however, leads to a gradual decrease in non-Gaussianity.
3.4 Impact of large optical depth on Skewness
We now shift our focus to the impact of high Hi 21-cm optical depth on $\delta T_{\rm b}$ and its statistics. If we choose a model where heating is inefficient as in $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$, $\tau$ would be sufficiently large throughout the entire CD as the spin temperature remains very low. As a result, $\delta T_{\rm b}$ calculated using exact equation is considerably suppressed compared to $\delta T_{\rm b}$ calculated using the approximated equation (see lower panel of fig.1). Further, the skewness of $\delta T_{\rm b}$ is suppressed compared to that estimated using the approximated $\delta T_{\rm b}$.
During the initial stages of CD, the skewness of the approximated and exact $\delta T_{\rm b}$ takes large negative values. This is because $\rm{Ly{\alpha}}$ radiation couples the Hi 21-cm spin temperature to the IGM kinetic temperature within regions surrounded by the sources, whereas the background IGM remains coupled with the CMBR temperature. This makes the Hi 21-cm signal negative inside the sources and consequently the overall signal negatively skewed. We do not see any significant difference between the exact and approximated $\delta T_{\rm b}$, mainly because Hi 21-cm optical depth is small during the initial stages of CD.
However, in Fig. 4 we see that the magnitude of skewness decreases as we go from redshift $z=18$ to lower redshifts. During this period, the $\rm{Ly{\alpha}}$ coupling percolates through the IGM and eventually results into a uniform $\rm{Ly{\alpha}}$ coupled IGM. The average spin temperature decreases and $\tau$ increases until the redshift around which the heating starts. This makes the higher-order terms of $\tau$ in equation (1) important. Consequently, the difference in skewness between the approximated and the exact $\delta T_{\rm b}$ also increases with decreasing redshift. This change can go up to $\sim 100\%$ in some cases considered here.
Once X-ray heating starts, kinetic and spin temperatures start increasing with decreasing redshift. Consequently, the optical depth starts to decrease with redshift. This makes the contribution of higher-order terms less important. Therefore, the exact $\delta T_{\rm b}$ is very similar to the approximated one toward the end of CD.
Due to very inefficient heating in models corresponding to $f_{\rm x}=0.1$, the IGM, including the $\rm{Ly{\alpha}}$ coupled regions near the sources, remains cold for an extended period of time. Therefore, the average $\delta T_{\rm b}$ remains negative and the PDF of $\delta T_{\rm b}$ remains negatively skewed up to redshift $z\sim 11$. Cold IGM in these models also lead to a larger optical depth making higher-order terms more impactful on the global signal as well as on the skewness. We see that (first two top-right panels in Fig. 4) the differences in skewness go up to $\sim 500\%$ and $\sim 75\%$ for models with $(f_{\rm x},M_{\rm h,min})=(0.1,10^{9}\,{\rm M}_{\odot})$ and $(0.1,10^{10}\,{\rm M}_{\odot})$, respectively.
X-ray heating is moderate in the model with $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. Therefore, the change in the skewness between approximated and exact $\delta T_{\rm b}$ goes up to $\sim 40\%$ at redshift $z\sim 14$ when the average optical depth becomes maximum in the model. Similarly, we see that the change is only up to $\approx 25\%$ at redshift $z\approx 12$ for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ for the same X-ray efficiency parameter. Models with $f_{\rm x}=1000$ exhibit similar features, but these appear at relatively earlier redshifts. Here, the maximum difference in skewness between the exact and approximated models is $\approx 60\%$ and $\approx 20\%$ respectively for $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $\rm 10^{10}\,{\rm M}_{\odot}$.
3.5 Redshift evolution of bispectrum
Bispectrum is the lowest order many-point statistics that is sensitive to non-Gaussianity. Bispectrum of a $\delta T_{\rm b}$ field within a volume $V$ can be written as,
$\delta^{\rm K}_{3\rm D}(\textbf{{k}}_{1}+\textbf{{k}}_{2}+\textbf{{k}}_{3})~{}B(\textbf{{k}}_{1},\textbf{{k}}_{2},\textbf{{k}}_{3})=V^{-1}\langle\Delta_{\rm b}(\textbf{{k}}_{1})\Delta_{\rm b}(\textbf{{k}}_{2})\Delta_{\rm b}(\textbf{{k}}_{3})\rangle,$
[…]
$k_{1}\geq k_{2}\geq k_{3};\quad{\rm}0.5\leq t,\,\mu\leq 1\quad{\rm and}\quad 2\mu t\geq 1.$
We have divided the $k$-space into $20$ logarithmic spherical shells for bispectrum estimation. Each combination of three shells forms a particular bin in the triangle shape and size space $(k_{1},\mu,t)$. The average size of the bin corresponds to $k_{1}$. Sampling of ($\mu,t$) space for a fixed $k_{1}$ represents different triangle shape configurations. We suggest the reader to Bharadwaj et al. (2020) and Shaw et al. (2021) for the details about the triangle shape and the bispectrum framework. In this work, we consider only the equilateral ($k_{1}=k_{2}=k_{3}$) and the squeezed limit triangles ($\textbf{{k}}_{1}\approx-\textbf{{k}}_{2}$, $\textbf{{k}}_{3}\approx 0$) for our bispectrum analysis.
Figure 5: This shows the redshift evolution of the spherically averaged dimensionless bispectrum for approximated and exact $\delta T_{\rm b}$, along with the percentage change between them, at the scale $k_{1}=0.16\,{\rm Mpc}^{-1}$. In the left panel, the blue solid and dashed lines represent bispectrum of approximated and exact $\delta T_{\rm b}$, respectively, for $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ while, the red solid and dashed lines represent the same for $M_{\rm h,min}=10^{10}{\rm M}_{\odot}$. In the right panel, the blue and red solid lines represent the percentage difference between the approximated and exact $\delta T_{\rm b}$ for the models with $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$, respectively.
Figure 6: Same as Fig. 5 for the scale $k_{1}=0.8\,{\rm Mpc}^{-1}$.
In Figs. 5 and 6, we show the redshift evolution of bispectra for both the exact and approximated $\delta T_{\rm b}$ at two different length scales. Fig. 5 shows results for a large length scale (small $k$) corresponding to $k_{1}=0.16$ Mpc-1 and Fig. 6 shows results for a small length scale corresponding to $k_{1}=0.8$ Mpc-1. We plot the spherically averaged dimensionless bispectrum $\Delta^{3}(k_{1},\mu,t)={k_{1}}^{3}{k_{2}}^{3}B(k_{1},k_{2},k_{3})/(2\pi^{2})^{2}$ (Majumdar et al., 2020) for both the equilateral and squeezed limit triangles. The bispectrum corresponding to the equilateral triangles finds the correlation between equal length scales and that corresponds to the squeezed limit configuration calculates the correlation between two equal length scales and one very large length scale, respectively.
In Fig. 5, we see that $\Delta^{3}(k_{1},\mu,t)$ is initially negative for $f_{\rm x}=0.1$ and $46.4$. At the beginning of CD, the spin temperature $T_{S}$ remains coupled to the CMBR temperature $T_{\gamma}$. This makes the Hi 21-cm signal nearly undetectable. However, as the first sources emerge, $\rm{Ly{\alpha}}$ photons from these sources begin coupling the spin temperature with the IGM kinetic temperature. As IGM kinetic temperature is smaller than the CMBR temperature, $\rm{Ly{\alpha}}$ coupled regions around the first sources appear as regions of negative $\delta T_{\rm b}$ against a zero background signal. This results in a negative bispectrum that continues as long as the $\rm{Ly{\alpha}}$ coupled regions remain isolated. However, these regions expand and percolate through the IGM. Eventually, the entire IGM spin temperature gets coupled to the IGM kinetic temperature. This widespread $\rm{Ly{\alpha}}$ coupling ultimately reduces the non-Gaussianity of the Hi 21-cm signal. We find that the bispectrum is negative (Fig. 5) from the start of CD until the redshifts $z\approx 14$ and $\approx 12$ for $M_{\rm h,min}=10^{9}$ and $10^{10}\,{\rm M}_{\odot}$, respectively, both for the equilateral and squeezed triangle configurations and $f_{\rm x}=0.1$. For a fixed $f_{\rm x}$ value, the overall nature in the features of bispectrum for both halo mass scenarios is almost same. The only difference is that, for higher $M_{\rm h,min}$ value, all the processes start at lower redshift. The redshifts until which the bispectrum remains negative due to $\rm{Ly{\alpha}}$ coupled regions is higher for higher $f_{\rm x}$, as can be seen in Fig. 5. However, we note that the bispectrum is positive since the beginning of CD for $f_{\rm x}=1000$ and the equilateral triangle configuration.
As the $\rm{Ly{\alpha}}$ coupling approaches completion, X-ray radiation from the first sources begins to heat the surrounding IGM. This heating raises both the kinetic temperature and the Hi 21-cm spin temperature of the IGM around these sources above the CMBR temperature. This makes $\delta T_{\rm b}$ positive inside heated regions, although the background IGM remains colder. The presence of large heated regions around the sources, buried in a colder IGM, makes the bispectrum positive. We see, in Fig. 5, that the bispectrum becomes positive at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, corresponding to $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively. The transition redshifts from negative to positive bispectrum are $z\approx 16$ and $\approx 14$ for $f_{\rm x}=46.4$ and $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively, in the equilateral triangle configuration. The transition occurs slightly later, at $z\approx 15$ and $\approx 13$, for the squeezed triangle configuration with the same halo mass thresholds. For $f_{\rm x}=1000$, X-ray heating is highly efficient, and the bispectrum remains positive from the beginning of CD in the equilateral triangle configuration. However for the squeezed triangle configuration, the bispectrum is initially negative for a short period of time, and becomes positive as X-ray heating starts. We see in this study that, for squeezed limit configuration, all the features appear at later stages compared to the equilateral one. The reason is that, in the squeezed-limit triangle configuration, there is a correlation involving one very large length scale, and astrophysical processes begin to affect such large scales at later times, i.e., at lower redshifts.
As the X-ray heating progresses, it gradually percolates through the IGM. This raises the entire IGM temperature above the CMBR temperature. The widespread heating again reduces the bispectrum. We see this effect in Fig. 5 for $f_{\rm x}=46.4$, for both equilateral and squeezed triangle configurations. Toward the final stage of heating, almost the entire IGM becomes heated, except a few underdense regions that are located far from the sources. This creates a scenario where $\delta T_{\rm b}$ is positive in most regions of the IGM but remains negative in those cold and underdense isolated pockets. This results in a large-scale negative bispectrum. For example, for $f_{\rm x}=46.4$ and $1000$, we see that the bispectrum becomes negative again at later times.
For $f_{\rm x}=1000$, X-ray heating is highly efficient. We find that the bispectrum becomes positive from the beginning of CD in the equilateral triangle configuration. Later, it goes negative at $z\approx 12.5$ and $\approx 10.5$ for $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively. However, for the squeezed triangle configuration, the bispectrum is initially negative and then becomes positive as the X-ray heating starts. Finally, it returns to negative value at relatively higher redshifts at $z\approx 15$ and $\approx 13$ for $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively, for the same reason as stated above.
Figure 6 shows the bispectra for $k_{1}=0.8\,{\rm Mpc}^{-1}$. The overall trends for $f_{\rm x}=0.1$ and $46.4$ are similar to those seen for $k_{1}=0.16\,{\rm Mpc}^{-1}$ (Fig. 5). However, there are some notable differences. For the squeezed triangle configuration and $f_{\rm x}=46.4$, the transition from a negative to a positive bispectrum happens much later—around redshift $z\sim 10.5$. Another distinct feature at $k_{1}=0.8\,{\rm Mpc}^{-1}$ is seen for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, $f_{\rm x}=1000$, and the equilateral triangle configuration. Here, the bispectrum remains negative throughout the entire CD period. Finally, for the squeezed triangle configuration with $f_{\rm x}=1000$, the transition from negative to positive bispectrum also occurs much later compared to the case of $k_{1}=0.16\,{\rm Mpc}^{-1}$.
3.6 Impact of large optical depth on bispectrum
The right columns of Figures 5 and 6 show the corresponding percentage change in the exact bispectra compared to the approximated bispectra calculated as, $\frac{\Delta^{3}_{\rm approx}-\Delta^{3}_{\rm exact}}{\Delta^{3}_{\rm exact}}\times 100\%$ as a function of redshift. We observe that, for most of the CD period, the magnitude of the exact bispectrum is suppressed compared to the approximate one. This suppression is mainly caused by the term $-\tau^{2}/2$ in the expression for the differential brightness temperature (equation 5), which reduces the value of $\delta T_{\rm b}$.
Since $\tau^{2}\propto\rho^{2}_{\rm HI}/T^{2}_{S}$, the suppression is not uniform across the field. It is stronger in regions with a higher neutral hydrogen density and lower spin temperature. As a result, the $\tau^{2}/2$ term affects the non-Gaussian signal (i.e., bispectrum and skewness) more than the global or Gaussian component of the signal. For example, the percentage change in the bispectrum is found to be $\sim 10\%$ at redshift $z=18$ across all models and triangle configurations. As redshift decreases, this change increases, reaches a maximum at an intermediate redshift, and then declines again. This trend can be explained if we look at the redshift evolution of Hi 21-cm optical depth. As shown in Fig.1, the average optical depth increases, peaks at a certain redshift, and then decreases at lower redshifts. The percentage difference in the bispectrum becomes largest when the optical depth is highest, consistent with what is observed in the right columns of Figs. 5 and 6. We find that the percentage change in the bispectrum increases up to $\sim 250\%$ at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, for both the equilateral and squeezed triangles at $k_{1}=0.16\,{\rm Mpc}^{-1}$. A similar trend is observed on smaller scales, e.g., at $k_{1}=0.8\,{\rm Mpc}^{-1}$ (see Fig. 6) where the change reaches up to $\sim 300\%$. This large change is due to the fact that both the IGM kinetic temperature and hence the spin temperature remain very low in this case because heating is highly inefficient. In models with moderate heating ($f_{\rm x}=46.4$), the percentage change is smaller, typically lies between $\sim 10\%$ to $\sim 200\%$. This is because heating is more efficient than in the low $f_{\rm x}$ case. This results in higher IGM and spin temperatures. As a result, the Hi optical depth becomes lower which reduces the impact of higher-order terms in $\tau$. For models with very efficient heating ($f_{\rm x}=1000$), the IGM and spin temperatures are large from the beginning of the CD. This makes the contribution of higher-order terms less significant. Consequently, the change in the bispectrum is small, around $10\%$, in these cases. However, we find that the change reaches up to $\sim 200\%$ at a few redshifts. We also find that the percentage change can sometimes be negative. This typically occurs when the bispectrum transitions from negative to positive or vice versa.
4 Impact on power spectrum
Figure 7: This shows the redshift evolution of the spherically averaged dimensionless power spectrum for the approximated and exact $\delta T_{b}$, along with the percentage difference between them, at scales $\rm k=0.16\ Mpc^{-1}$ (upper two rows) and $\rm k=0.8\ Mpc^{-1}$ (lower two rows). In the first and third rows, the blue solid and dashed lines represent the power spectrum of the approximated and exact $\delta T_{b}$, respectively, for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ model, whereas the red solid and dashed lines represent the same for the $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ model. In the second and fourth rows, the blue and red solid lines show the corrsponding percentage differences between the approximated and exact cases for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ models, respectively.
Although our main goal is to investigate the impact of higher-order terms of $\tau$ on the non-Gaussian features of the 21-cm signal from the CD, it is also worthwhile to study the Hi power spectrum in this context. Measuring the power spectrum is the primary objective of ongoing and upcoming CD and EoR radio interferometric experiments. Significant progress has already been made in modelling the power spectrum (Furlanetto et al., 2004; Mesinger & Furlanetto, 2007; Datta et al., 2012; Jensen et al., 2013), placing upper limits using ongoing experiments (Acharya et al., 2024; Kolopanis et al., 2019; Trott et al., 2020; Abdurashidova et al., 2023), and interpreting these measurements (Mondal et al., 2020; Abdurashidova et al., 2022; Greig et al., 2021; Ghara et al., 2025).
In Fig. 7, we show the redshift evolution of spherically averaged power spectra of both approximated and exact $\delta T_{\rm b}$ at the length scales $k=0.16$ Mpc-1 and $k=0.8$ Mpc-1. We can see that, the power spectrum initially increases with decrease in redshift, and then decreases beyond a certain redshift. It starts decreasing earlier for higher $f_{\rm x}$ value. However, unlike bispectrum evolution which is a marker of different astrophysical events through its sign flip, power spectrum is positive throughout the entire CD period.
We can see, for $k=0.16$ Mpc-1, that at the beginning of the CD power spectrum increases as the $\rm{Ly{\alpha}}$ radiation starts coupling the spin temperature with the kinetic temperature. However, as the average $\tau$ is small at the beginning of CD, the change in the power spectrum between approximated and exact cases is small. As time progresses, $\rm{Ly{\alpha}}$ coupling becomes a dominating astrophysical event and as $\rm{Ly{\alpha}}$ radiation spreads throughout the IGM, the average $\tau$ becomes maximum around that redshift. Therefore, the change in the power spectrum is also maximum at that redshift. After that, when X-ray radiation starts heating the IGM, the average $\tau$ starts decreasing, which also decreases the change in the power spectrum. One notable difference here is that power spectra do not show any sign flip which can discriminate between these two different astrophysical events, whereas the bispectrum evolution shows a sign flip between these events (see Fig.5). Finally, when IGM gets almost entirely heated the power spectrum and its change both decreases.
Similar behaviour is seen in all other scenarios. Considering the model with $f_{\rm x}=46.4$ and $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ for $k_{1}=0.16\,{\rm Mpc}^{-1}$, we see the percentage change between approximated and exact power spectra increases with decrease in redshift as average $\tau$ increases while $\rm{Ly{\alpha}}$ coupling dominates over X-ray heating. The maximum change is $\approx 20\%$ around $z=15$ for this model. For higher halo mass threshold, such as $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ with same $f_{\rm x}$ value we see similar changes but at later redshifts. For models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ the change goes up to $\approx 30\%$ at redshift $z\approx 11.5$ and $z\approx 10.5$, respectively. The change reaches up to merely $\approx 8\%$ at redshift $z\approx 16.5$ and $z\approx 14$ for models with $f_{\rm x}=1000$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=1000$, $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, respectively.
Similar qualitative behaviour is found at smaller length scales. The percentage changes in the power spectrum at $k=0.8$ Mpc-1 are $\approx 28\%$ and $\approx 30\%$ for models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, respectively. For higher $f_{\rm x}$ values, the changes are similar to that of $k=0.16\,{\rm Mpc}^{-1}$. For example, changes for $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ models are up to $\approx 20\%$ and for $f_{\rm x}=1000$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=1000$, $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ models the changes are as high as $\approx 8\%$ and $\approx 11\%$ at $z\approx 16.5$ and $z\approx 13.5$, respectively.
5 Importance of different Higher-Order Terms of $\tau$ in Bispectrum Estimation
Figure 8: This shows the percentage differences between the exact and approximated bispectra of $\delta T_{\rm b}$ for the model with $f_{\rm x}=0.1$ and $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$, at redshifts $z=12.3$, $14$, and $15.6$. The approximated $\delta T_{\rm b}$ includes contributions up to the first (blue), second (green), and third order (red) terms of $\tau$. The left and right panels correspond to the equilateral triangle configuration and the squeezed limit, respectively.
It would be interesting to investigate how many higher-order terms of $\tau$ need to be retained in equation (5) to accurately model $\delta T_{\rm b}$, especially for capturing the non-Gaussian features of the signal. Furthermore, it is important to understand how each higher-order term of $\tau$ contributes to the bispectrum. Therefore, expanding the exact expression of $\delta T_{\rm b}$ (1) in powers of $\tau$ and determining which orders are important would be a useful step for accurate modelling and analysis.
In Fig. 8, we show the percentage difference between the exact bispectrum and those estimated using up to the first, second, and third order terms of $(1-e^{-\tau})$ for $f_{\rm x}=0.1$ and $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. We see that, the difference is maximum, reaching up to a few $100\%$, when using only up to the first-order term for all scenarios considered. However, the difference reduces significantly to $\sim 10\%$ after including the second-order term. When terms up to the third order are included in estimating $\delta T_{\rm b}$, the differences become negligibly small, typically $\sim 1\%$ or less. We have shown the result for three different redshifts $12.3$, $14$ and $15.6$. We choose this range of redshifts as average $\tau$ is relatively high in this range, leading to a large difference in the percentage change. However, this trend holds true for all values of $k_{1}$, redshifts, models, and triangle configurations considered in this study.
6 Summary and Discussions
The Hi 21-cm optical depth, which is proportional to the neutral hydrogen density and inversely proportional to the spin temperature, can become quite large during the Cosmic Dawn (CD). This is because the spin temperature is expected to remain very low before X-ray heating begins. Our simulations show that, in some CD models, the optical depth can reach values as high as $\approx 0.55$, with nearly $50\%$ of the simulation pixels showing optical depths greater than $0.1$. Therefore, the commonly used approximation that the optical depth is much less than one ( $\tau<<1$) which simplifies the expression $[1-\exp(-\tau)]\approx\tau$ may not hold during this epoch.
To revisit the impact of this assumption on $\delta T_{b}$ and its associated statistics, we performed a suite of numerical simulations spanning a wide range of astrophysical parameters, including the X-ray heating efficiency ($f_{\rm x}$) and the minimum halo mass ($M_{\rm h,min}$). These simulations model $\rm{Ly{\alpha}}$ coupling, X-ray heating, spin temperature evolution, and ionization using a 1D radiative transfer code, named grizzly. We used them to study how large optical depths affect the statistical properties of the 21-cm signal, including the power spectrum, skewness, and bispectrum of the differential brightness temperature $\delta T_{\rm b}$.
We find that $\delta T_{\rm b}$ is significantly suppressed when the full expression with $[1-\exp(-\tau)]$ is used, particularly in regions with high neutral hydrogen density and low spin temperature. This suppression is not uniform across the volume but is stronger in regions where the ratio $\rho_{\rm HI}/T_{s}$ is higher. As a result, a new kind of non-Gaussianity (due to higher order terms of $\tau$ ) emerges with a sign opposite to that of the existing bispectrum. In the model with $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=0.1$, the skewness changes by up to $\sim 500\%$. For the same halo mass, the change is about $40\%$ and $60\%$ for $f_{\rm x}=46.4$ and $f_{\rm x}=1000$, respectively. For a higher halo mass threshold of $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, the skewness changes by approximately $75\%$, $25\%$, and $20\%$ for $f_{\rm x}=0.1$, $46.4$, and $1000$, respectively.
We also studied the impact on the bispectrum in both equilateral and squeezed triangle configurations across all redshifts during CD. We analyzed two different scales, $k_{1}=0.16\,{\rm Mpc}^{-1}$ and $0.8\,{\rm Mpc}^{-1}$, representing large and small scales. For low-heating models ($f_{\rm x}=0.1$) with both $M_{\rm h,min}=10^{9},{\rm M}_{\odot}$ and $10^{10},{\rm M}_{\odot}$, where the optical depth remains high throughout the CD, the change in the bispectrum is consistently large—reaching up to $\sim 300\%$ at both scales and triangle configurations. However, in models with moderate and high X-ray heating efficiency ( $f_{\rm x}=46.4$ and $1000$), the changes are smaller—typically around $\sim 10\%$ to $\sim 200\%$, regardless of halo mass or triangle shape.
Further, we investigated how many terms in the $\tau$-expansion are needed to accurately model $\delta T_{\rm b}$, especially to capture its non-Gaussian features. We find that including terms up to third order in $\tau$ reduces the difference to less than $\sim 1\%$. This result holds across all $k$-values, redshifts, models, and triangle configurations considered in our study.
Finally, we note that our entire analysis has been carried out in real space. Including the effect of peculiar velocities would further enhance the optical depth. This could amplify the impact on the 21-cm signal.
Data Availability
The data underlying this work will be shared upon reasonable request to the corresponding author.
7 Acknowledgments
IN acknowledges financial support from the Swami Vivekananda Merit-cum-Means Scholarship (SVMCM), Govt. of West Bengal, India. IN also acknowledges support from the DST-WISE Fellowship (DST/WISE-PhD/PM/2023/104), Govt. of India. KKD acknowledges financial support from ANRF (Govt. of India) through a SERB MATRICS scheme (MTR/2021/000384). The authors also acknowledge the use of computational resources provided by the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune. RG acknowledges support from SERB, DST Ramanujan Fellowship no. RJF/2022/000141.